Prev Next

SW&IT Meeting Minutes 2022-12-12

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20221128
LR20221128 - Bec, Matthieu 80.0
LR20221128 - Mechtley, Brandon 72.0
LR20221128 - Eychaner, Glenn 80.0
LR20221128 - Edmeades, David 80.0
LR20221128 - Grenz, Paul 36.0
LR20221128 - Hooper, Stephen 0.0
LR20221128 - Zhang, Yang 80.0
LR20221128 - Zhang, Xianyu 96.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, ATT, Management
    • tcs sim VM
    • test wfs-srv
    • SOUL team visit

  • This week
    • career conversations
    • tcs sim VM
    • explore swgdev04-alma9.tucson.lbto.org
    • obs station iteration, 5x in 507
    • look at lib-telemetry and h5 patch (needed or not)

YZ

  • Previous
    • TMS support work and tests (Tuesday night plan cancelled at the last minute due to bad weather, Thursday night on-sky test cancelled due to weather, but instead we conducted half-night closed dome test).
    • Working on various archive utilities and tools.
    • Supplied instruments people the proper ways to start/stop MODS alarms, also investigated some minor MODS utility issues.
    • All staff meeting on Friday.
    • Submitted career conversation in UAccess.
  • Next:
    • May support further TMS tests or data analysis.
    • Revise TMS wiki documentation and fixing problems in TMS code as discussed during the tests.
    • Continue to work on archive utilities.
    • Vacation until Jan 09.

DE

  • Last week
    • UITS ticket for KVM connectivity; simple VLAN swap.
    • Ordered parts for AO server upgrades
    • User desktop sundries
  • This week
    • Office trip
      • Prep to move luci-new into server room; make sure we have a switch port on VLAN 105
      • Verify KVM dongle situation, but planning to redeploy one
      • JP laptop get and re-OS
    • Look into virtualizing tcs-test as discussed

GCE

Last week
  • Deployed LUCI 2.15.10 with a few minor fixes, including warnings when ReentrantLock.tryLock() fails. Also deployed new test GUIs.
  • IT 8498 (Stale LUCI FITS header info): Provided a document detailing where the info for the LUCI FITS header keywords is acquired. Also figured out why stale info is being propagated from one observing script to the next: CommitManager.updateObservationDescriptor() copies the data each commit.
  • IT 8774 (Non-sidereal tracking FITS headers): Implemented in LUCI code; to be deployed next week.
  • More discussion of the TCS simulator and swing arms.
  • Committed to creating a kickstart script while I am installing software on the new LUCI computer.
This week
  • Fix the observation GUI so that it clears the ObservationDescriptor data each time a new script is loaded, and on exit.
  • Install LUCI software onto the new LUCI computer.
  • More Observer GUI improvements.

BM

  • Last week
    • spare-soul-wfs
    • it 8588 camera lens temp warning
    • Friday: misc patch, hk+rr/misc testing
  • This week
    • finish comments for s-lbti-sx acceptance
    • career conversation
    • it 8762 (rerotator)
    • finish up kickstart script
    • Xianyu's HK+rr patch went OK, but still needs night time checkout (next DEC 24th ... could leave instructions or wait until after NY). Will see if we can do another daytime test for "misc" patch.

PSG

  • Last week
    • OSS ARGOS sim update (no MCSPU sim)
    • PMAC program merge
  • This week
    • Finish ARGOS sim
    • Finish PMAC merge

SH

  • to1 issues - nothing overt, but stability is waning
  • WiFi network on the mountain. No move on level 5.
  • build system IRTC, need to check it in
  • ARGOS computer stuff -- various hardware and software problems.
  • DNS things

next:

  • ITGC
  • Move Gene's files.
Topic revision: r9 - 12 Dec 2022, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback